OpenStudy (moongazer):

6 people can paint 4 walls in 43 minutes. How many minutes will it take for 7 people to paint 7 walls? Round to the nearest minute.

OpenStudy (anonymous):

7 ppl can paint 7 walls in 43*4*6/49=35 min

OpenStudy (moongazer):

could you show the solution? I already have the answer :)

OpenStudy (moongazer):

can anyone explain it to me clearly?

OpenStudy (moongazer):

@vamgadu your answer is wrong

OpenStudy (anonymous):

hey 1 person to paint one wall takes 43*6*4 mins so 7persons to paint 7 walls take 43*6*4/49 mins remember number of people and amount of work is inversely prop to time.

OpenStudy (moongazer):

the answer should be 65

OpenStudy (anonymous):

oh wait i am wrong

OpenStudy (anonymous):

ok so 6 people can paint 4 walls in 43 minutes. this means that 6 people can paint 1 wall in (43/4) minutes this would be 6 times faster than 1 person painting 1 wall, so we multiply by 6 one person painting one wall: 6*(43/4) now we need to divide this by seven, as seven people will be seven times as fast: (6*(43/4) )/7 now we multiply by seven because we're painting 7 walls: 6*(43/4) =64.5 65 to the nearest minute

OpenStudy (anonymous):

eigenschmeigen is correct

OpenStudy (moongazer):

It still confuses me a little bit :)

OpenStudy (anonymous):

where do you get confused? which bit starts to confuse you? i'll try and explain more

OpenStudy (moongazer):

why is it that 6 people can paint 1 wall in (43/4) minutes?

OpenStudy (anonymous):

First understand that the problem presented means that it takes 6*43=258 minutes of man-work to paint 4 walls Which means that it takes 258/4=64.5 minutes of man-work to paint each wall. Therefore, it takes 64.5*7=451.5 of man-work to paint 7 walls. Finally we derive the answer to the problem as so: it takes 451.5/7=64.5 minutes for 7 people to paint 7 walls. i asked it on this site. http://www.algebra.com/tutors/students/your-answer.mpl?question=607974 it helps when open study is going slow. Hope it helped!